# Thermoelectric energy conversion in molecular junctions out of equilibrium

### v1

@@ Abstract
Understanding time-resolved quantum transport is crucial for developing next-generation quantum technologies, particularly in nano- and molecular junctions subjected to time-dependent perturbations. Traditional steady-state approaches to quantum transport are not designed to capture the transient dynamics necessary for controlling electronic behavior at ultrafast time scales. In this work, we present a non-equilibrium Green’s function formalism, within the recently-developed iterated generalized Kadanoff-Baym ansatz ($i$GKBA), to study thermoelectric quantum transport beyond the wide-band limit approximation (WBLA). We employ the Meir-Wingreen formula for both charge and energy currents and analyze the transition from Lorentzian line-width functions to the WBLA, identifying unphysical divergences in the latter. Our results highlight the importance of finite-bandwidth effects and demonstrate the efficiency of the $i$GKBA approach in modeling time-resolved thermoelectric transport, also providing benchmark comparisons against the full Kadanoff-Baym theory. We exemplify the developed theory in the calculation of time-resolved thermopower and thermoelectric energy conversion efficiency in a cyclobutadiene molecular junction.
@@ I Introduction
While thermoelectric devices are traditionally macroscopic and valued for their stable, time-independent performance, a new generation of applications demands precise control of charge and heat transport at atomic and molecular scales, as well as ultrafast time scales. In these regimes, conventional steady-state descriptions break down, and the interplay between energy conversion and quantum coherence must be addressed. Recent studies have explored these challenges in a variety of systems, including ferromagnetic-superconducting interfaces [1], condensed-matter platforms [2], and qubit-based thermal devices [3]. Coherence effects in mesoscopic systems enable applications such as precision thermometry and efficient refrigeration [4], while thermoelectric phenomena at the atomic and molecular scale highlight the interplay between quantum heat flow, electronic transport, and non-equilibrium thermodynamics [5]. Thermoelectric properties of single-molecule junctions have been explored both experimentally and theoretically [6], with a focus on inelastic transport effects and the Seebeck coefficient [7], and the optimization of the thermoelectric figure of merit [8, 9]. Thermoelectric efficiency and energy harvesting have been investigated in multi-terminal quantum-dot systems operating as thermal engines [10, 11, 12, 13, 14, 15]. These advances underline the importance of thermoelectric transport in enabling quantum technologies, energy-efficient devices, and thermal control techniques at the nanoscale.
@@
The study of time-resolved quantum transport is essential for understanding the behavior of electrons in nano- and molecular junctions under external driving mechanisms, such as time-varying voltages, electromagnetic fields, or sudden changes in system parameters [16, 17, 18, 19]. With miniaturization to atomic and molecular scales, quantum devices push their operating frequencies into the terahertz range, reaching electronic time scales on the order of femtoseconds [20, 21]. Hybrid optoelectronic platforms built from low-dimensional materials (e.g., graphene and topological insulators) exploit high carrier mobility and strong light-matter coupling to function in this regime [22] and have demonstrated active control of electromagnetic signals at sub-picosecond scales using current fabrication techniques [23]. This branch of research is especially relevant for the design of next-generation quantum materials and technologies, where controlling electronic behavior at ultrafast time scales is critical [24, 25, 26, 27, 28]. Traditional approaches to quantum transport, relying on linear response theory and steady-state assumptions [29, 30, 31], are not designed to capture the complex dynamics of systems under strong and time-dependent perturbations. Alternative approaches such as renormalization group and quantum Monte Carlo methods are typically perturbative in the system-lead coupling and become computationally prohibitive in higher dimensions [32]. Time-dependent density-functional theory relies on approximate exchange-correlation functionals, which often fail to capture non-local correlations [33, 34, 35], while scattering wave function methods are generally limited to weakly interacting systems [36].
The non-equilibrium Green’s function (NEGF) formalism serves as a powerful framework to study time-dependent processes [37, 38, 39]. By resolving the time evolution of many-body quantum systems via the Kadanoff-Baym equations (KBE), the NEGF formalism allows for tracking the transient behavior of currents, charge distributions, and other observables in response to sudden changes [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50]. This time-domain analysis is particularly useful for understanding the immediate response of quantum systems to external drives, as well as for studying relaxation dynamics as the system evolves toward the stationary state [51]. Within the NEGF theory, a compelling technique for time-resolved transport is the generalized Kadanoff-Baym ansatz (GKBA) [52, 53], especially in its recently-developed time-linear framework [54, 55, 56, 57, 58, 59]. While the previous works have established and applied a time-dependent Meir-Wingreen formula for the charge current [59, 60], the corresponding formula for the energy current and thermal transport has only very recently been discussed within the time-linear GKBA formulation [61].
@@
The coupling between a quantum system, such as a molecular junction or quantum dot, and its surrounding leads or electrodes (see Fig. 1) is commonly described by a tunneling rate or line-width function. It characterizes how the states in the central region (e.g., the molecular junction) are broadened due to their interaction with the continuum of states in the leads and determines the rate at which electrons tunnel between the quantum system and the leads. The bandstructure or density of states (DOS) of the leads is decisive in shaping the energy dependence of the line-width function. Due to a considerable computational simplification, a common approach is the so-called wide-band limit approximation (WBLA) where the lead DOS is assumed flat or featureless [62, 63]. However, in a more realistic transport setup, the line-width function deviates from the WBLA form, i.e. having some energy dependence, directly influencing the overall transport characteristics.
@@
Therefore, treating time dependence, interactions, and energy-dependent lead coupling is crucial for the accurate modeling of thermoelectric quantum transport in driven nanoscale systems. In this work, we develop a non-equilibrium Green’s function framework, expanding on our recent Letter [61], to address these topics on an equal footing. We present a detailed derivation of the iterated generalized Kadanoff-Baym ansatz ($i$GKBA) concentrating on a Lorentzian tunneling-rate function. We employ the Meir-Wingreen formula for both charge and energy currents and analyze the reduction from the Lorentzian to the WBLA description. The conventional GKBA approach, a subset of the $i$GKBA scheme, is found to be insufficient for studying energy currents, as it leads to an unphysical divergence within the WBLA. This issue is resolved by the finite-bandwidth $i$GKBA as demonstrated by extensive benchmark results and comparison with the full Kadanoff-Baym equation approach in a carbon-based molecular junction. Ultimately, we put the developed $i$GKBA scheme to the test by investigating the time-dependent thermoelectric energy conversion in a molecular junction under a temperature gradient.
@@ II.2 Non-equilibrium Green’s function theory
The central object of the non-equilibrium Green’s function theory, in describing the out-of-equilibrium quantum-transport setting of Eq. (1) is the one-electron lesser ($<$) Green’s function (GF) [38]
@@
[…]
@@
[…]
where $h_{mn}^{\mathrm{HF}}(t)=h_{mn}(t)+V_{mn}^{\mathrm{HF}}(t)$ is the one-electron Hamiltonian including the Hartree-Fock (HF) potential $V_{mn}^{\mathrm{HF}}(t)=\sum_{pq}(v_{mpqn}(t)-v_{mpnq}(t))\rho_{qp}(t)$. The self-energy $\Sigma=\Sigma_{c}+\Sigma_{\mathrm{em}}$ is composed of the correlation and embedding parts. The correlation self-energy can be specified at different levels of approximations: second-order Born [64, 65], $GW$ and $T$-matrix [66, 67], or based on the non-equilibrium dynamical mean-field theory [68, 69, 70]. The embedding self-energy accounts for the system being open to an environment consisting of the leads. The superscripts ‘$\mathrm{A}$’ and ‘$\mathrm{R}$’ denote the advanced and retarded components, respectively.
@@ III.1 Collision integrals
[…]
@@ III.2 Reconstruction equations
It is convenient to split the lesser GF into the retarded- and advanced-like ($\mathrm{R}$/$\mathrm{A}$) components
@@
[…]
@@
[…]
One obtaines the “standard” GKBA equations by neglecting the memory corrections (given by the integral over $y$). The iterated GKBA ($i$GKBA) is obtained by iterating the equations further, i.e., inserting the GKBA expression for $G^{<}_{\mathrm{A}}$ from the second equation into the first, and $G^{<}_{\mathrm{R}}$ from the first equation into the second. In order to efficiently deal with different kinds of convolution integrals that appear in this work, we introduce short-hand notations:
[…]
@@
This allows to rewrite the $i$GKBA equations as follows:
@@ III.3 Embedding self-energy components
[…]
@@
×$\qquad\times\mathcal{F}\left[\omega^{\nu}\Gamma_{\alpha}(\omega)\right](t-t^{\prime})\theta(t-t^{\prime}),$ (17a)
@@
×$\qquad\times\mathcal{F}\left[\omega^{\nu}f_{\alpha}(\omega)\Gamma_{\alpha}(\omega)\right](t-t^{\prime}),$ (17b)
@@
$\bar{\eta}_{\alpha\ell}$ =$=\begin{cases}i\frac{\gamma_{\alpha}}{2}\Omega_{\alpha}f_{\alpha}(\epsilon_{\alpha}-i\Omega_{\alpha})&\ell=0,\\ -\frac{\eta_{\ell}}{\beta_{\alpha}}\Gamma_{\alpha}\left(\mu_{\alpha}-i\frac{\zeta_{\ell}}{\beta_{\alpha}}\right)&\ell\geq 1;\end{cases}$ (23)
@@
$\bar{\mu}_{\alpha\ell}$ =$=\begin{cases}\epsilon_{\alpha}-i\Omega_{\alpha}&\ell=0,\\ \mu_{\alpha}-i\frac{\zeta_{\ell}}{\beta_{\alpha}}&\ell\geq 1.\end{cases}$ (24)
@@ III.4 Working scheme
Figure 2: Derivation of the EOMs for the $i$GKBA correlators. Correlators that originate from the self-energy terms in Eqs. (33) are framed in red. Correlators that appear already at the GKBA level are shaded.
@@
$-i\frac{d}{dt}G^{\mathrm{A}}(t^{\prime},t)=\delta(t^{\prime}-t)\\ +G^{\mathrm{A}}(t^{\prime},t)h^{\mathrm{HF}}(t)+\bigl{[}G^{\mathrm{A}}\cdot\Sigma^{\mathrm{A}}\bigr{]}(t^{\prime},t),$ (33b)
@@ IV Benchmark simulations for a quantum-dot system
There are different levels of approximations outlined in Sec. III. Inserting the GKBA iteratively back into the reconstruction equations (11) and (12) resulted in the full set of $14$ EOMs, which is referred to $i$GKBA. Neglecting the correction terms $\mathcal{A}$ in the collision integral (29) reduces to the GKBA framework (only shaded correlators in Fig. 2 are considered) with the embedding self-energies described beyond the WBLA, through the Lorentzian line-width function (16). This finite-bandwidth Lorentzian description can be applied systematically to both the collision integrals (29) and the equations of motion for the advanced propagators (33). We refer to this scheme as “GKBA ($\Omega,\ G^{\mathrm{A}}$)”. Alternatively, the retarded/advanced propagators can be approximated at the wide-band approximation level, i.e., the bandwidth $\Omega$ is finite for the embedding self-energies, but the equations of motion (33) include an effective Hamiltonian $h_{\mathrm{eff}}=h^{\mathrm{HF}}-i\gamma/2$ with the frequency-independent line-width matrix ($\gamma=\sum_{\alpha}\gamma_{\alpha}$; see Eq. (16)). We refer to this as “GKBA ($\Omega,\ h_{\mathrm{eff}}$)” with red-framed correlators in Fig. 2 being additionally neglected. The limiting case, $\Omega\to\infty$ is the wide-band limit approximation, simply referred to as “GKBA (WBLA)”: For the charge density and current, we refer to Ref. [59] and for the energy current, an alternative derivation is presented in the Supplemental Material [78]. The ($i$)GKBA calculations were performed using the cheers code [79, 80].
@@
We focus on a thermoelectric transport setup consisting of a two-terminal quantum dot with a temperature gradient and a time-dependent chemical potential difference. This model, studied analytically in Ref. [81], can capture complex dynamic thermoelectric behavior, and our benchmark calculations illustrate how different operating regimes impact efficiency, charge and energy-current responses of the system, and the validity range of the developed theoretical formulation in Sec. III.
@@
Figure 3: Time-dependent charge density $\rho$ (a), charge current $J_{L}$ (b), and energy current $J_{L}^{E}$ (c) at the left-lead interface with the quantum dot being weakly coupled to the leads, $\gamma=0.1$. Inset in panel (c) displays the equilibrium energy current, before the sudden gate voltage is switched on, in terms of the number of poles.
@@
We first fix the quantum dot energy level to $\epsilon_{0}(t)=0.5+u(t)$ with a sudden gate voltage $u(t)=2.5\theta(t)$. The chemical potentials of the left and right leads are first set to $\mu_{L}=\mu_{R}=0$ and there is no temperature gradient $\beta_{L}=\beta_{R}=10$. In all calculations, we set $\Omega_{\alpha}=\Omega$ for all leads $\alpha$. In Fig. 3, the time-dependent response is analyzed in terms of the charge density on the quantum dot $\rho$ and the charge and energy current at the left-lead interface $J_{L}$, $J_{L}^{E}$, when the quantum dot is weakly coupled to the leads $\gamma_{L}+\gamma_{R}=\gamma=0.1$. We note that $\gamma$ is a scalar for the single-level quantum-dot system. The time evolutions are started from an initially disconnected system $\rho(t\to-\infty)=0$ and adiabatically ramping the coupling between the quantum dot and the leads. As discussed in Ref. [61], with aligned chemical potentials and no temperature gradients, the equilibrium state should exhibit no currents. The comparison of density and charge current to the analytical result of Ref. [81] (“Kara Slimane et al.” in Fig. 3) is excellent already at the WBLA level. When the leads are described with a finite-width Lorentzian, GKBA ($\Omega=10,\ h_{\mathrm{eff}}$) and GKBA ($\Omega=10,\ G^{\mathrm{A}}$), there is a small change in the equilibrium density $\rho(t=0)$. This has no apparent effect on the charge current. The $i$GKBA results follow the same behavior.
@@
For the energy current, Fig. 3(c), the situation changes qualitatively. Although the result of Ref. [81] is obtained by numerical integration, the energy current still converges to a physically expected value at equilibrium $J_{L}^{E}(t=0)=0$. The $i$GKBA calculation matches this fairly well, although the finite-bandwidth $\Omega=10$ does not precisely correspond to the WBLA. In contrast, for the standard GKBA approach, at different levels of approximation, the corresponding result changes with the number of poles in the expansion (21). The main panels in Fig. 3 include $N_{p}=30$ poles, and the pole count is varied at the inset of panel Fig. 3(c), showing a slow divergence (logarithmic) in the case of WBLA. Interestingly, the GKBA calculations with finite-width Lorentzians do converge and the convergence is relatively fast, around $N_{p}\gtrsim 10$ poles, but they converge to an unphysical value ($\neq 0$). The overall mismatch in the energy currents is still fairly minimal at weak coupling.
@@
Figure 4: Same as Fig. 3 but the quantum dot being strongly coupled to the leads, $\gamma=1.0$.
@@
The $i$GKBA scheme accurately resolves the energy currents in the strong coupling regime as well; see Fig. 4. Here, the parameters are otherwise the same as in the weak-coupling calculation (Fig. 3) but the coupling strength is varied so that $\gamma=1.0$. Interestingly, the charge density and current [Fig. 4(a) and Fig. 4(b)] are accurately described by the standard GKBA, but the energy current again diverges for the WBLA case with increasing pole count. We clearly see that when the coupling strength is increased, the standard GKBA at the level of WBLA becomes entirely insufficient. The GKBA calculations with finite-band Lorentzian converge again around $N_{p}\gtrsim 10$ poles, but there is still a noticeable shift in the equilibrium energy current. The $i$GKBA corrects this unphysical behavior, but it is worth noting that the finite-bandwidth Lorentzian with $\Omega=10$ description deviates more from the result of Ref. [81], obtained at the WBLA, when the coupling is stronger. Still, it is required to go beyond GKBA to fix the problem of unphysical energy currents in equilibrium.

### v2

@@ Abstract
Understanding time-resolved quantum transport is crucial for developing next-generation quantum technologies, particularly in nano- and molecular junctions subjected to time-dependent perturbations. Traditional steady-state approaches to quantum transport are not designed to capture the transient dynamics necessary for controlling electronic behavior at ultrafast time scales. In this work, we present a non-equilibrium Green’s function formalism, within the recently-developed iterated generalized Kadanoff-Baym ansatz (iiGKBA), to study thermoelectric quantum transport beyond the wide-band limit approximation (WBLA). We employ the Meir-Wingreen formula for both charge and energy currents and analyze the transition from Lorentzian line-width functions to the WBLA, identifying unphysical divergences in the latter. Our results highlight the importance of finite-bandwidth effects and demonstrate the efficiency of the iiGKBA approach in modeling time-resolved thermoelectric transport, also providing benchmark comparisons against the full Kadanoff-Baym theory. We exemplify the developed theory in the calculation of time-resolved thermopower and thermoelectric energy conversion efficiency in a cyclobutadiene molecular junction.
@@ I Introduction
While thermoelectric devices are traditionally macroscopic and valued for their stable, time-independent performance, a new generation of applications demands precise control of charge and heat transport at atomic and molecular scales, as well as ultrafast time scales [1, 2, 3]. In these regimes, conventional steady-state descriptions break down, and the interplay between energy conversion and quantum coherence must be addressed.
@@
Recent studies have explored these challenges in a variety of systems, including ferromagnetic-superconducting interfaces [4], condensed-matter platforms [5], and qubit-based thermal devices [6]. Coherence effects in mesoscopic systems enable applications such as precision thermometry and efficient refrigeration [7]. Thermoelectric phenomena at the atomic and molecular scale highlight the interplay between quantum heat flow, electronic transport, and non-equilibrium thermodynamics [8]. Single-molecule junctions have been explored both experimentally and theoretically [9], with a focus on inelastic transport effects and the Seebeck coefficient [10], and the optimization of the thermoelectric figure of merit [11, 12, 13]. Thermoelectric efficiency and energy harvesting have been investigated in quantum-dot systems operating as thermal engines [14, 15, 16, 17, 18, 19]. These advances underline the importance of thermoelectric transport in enabling quantum technologies, energy-efficient devices, and thermal control at the nanoscale.
@@
In this work, we focus on the dynamical aspects of quantum transport, as manifested in the behavior of electrons in nano- and molecular junctions under external driving mechanisms such as time-varying voltages, electromagnetic fields, or sudden changes in system parameters [20, 21, 22, 23, 24]. With miniaturization to atomic and molecular scales, quantum devices reach operating frequencies in the terahertz range, corresponding to electronic time scales on the order of femtoseconds [1, 25]. Optoelectronic platforms based on low-dimensional materials exploit high carrier mobility and strong light-matter coupling in this regime [26] and have demonstrated active control of electromagnetic signals on sub-picosecond scales [2]. Controlling electronic behavior at ultrafast time scales [27, 28, 29, 30, 31] is therefore critical for designing next-generation quantum materials and technologies.
Theoretical approaches to time-dependent quantum transport can be broadly divided into two categories. One strategy is to propagate scattering states in time [32, 33], as done in microcanonical formalism [34, 35, 36], time-dependent density-functional theory [37, 38, 39] or in density-matrix renormalization group [40] techniques. These methods scale linearly with physical propagation time and can, in principle, capture strong correlations, but they require an explicit microscopic description of both the device and its reservoirs. This explicit treatment can become computationally expensive even for noninteracting systems [41] and poses conceptual challenges at finite temperature [42, 43].
The second strategy is based on non-equilibrium Green’s functions (NEGF) [44, 45, 46], where the dynamics are governed by the Kadanoff-Baym equations (KBE). They describe the time evolution of single-particle Green’s functions and have also found applications beyond transport, notably in ultrafast phenomena and time-resolved spectroscopy [47, 48, 49, 50, 51, 52, 53, 54, 55, 56]. Within this framework, transient currents can be computed using the time-dependent Meir-Wingreen [57, 20] or Landauer-Büttiker [58, 59, 60, 61, 62] formulas. However, direct numerical solutions of the KBE are limited by their cubic scaling with physical time, which has motivated the development of reconstruction schemes and related approximations [63, 64]. The most prominent example is the generalized Kadanoff-Baym ansatz (GKBA), which reduces the scaling to linear time [65, 66, 67, 68, 69, 70] and performs well for charge transport in the wide-band limit [70, 71]. However, these approximations discard essential spectral information and fail for energy currents.
Our work overcomes this bottleneck by developing the iterated GKBA (iiGKBA), a time-linear NEGF method that remains well defined for energy currents beyond the wide-band limit and is well suited to describe energy-dependent quantities such as heat currents. This study builds on our recent Letter [72], where we employed a time-dependent Meir-Wingreen formula for energy currents and thermal transport. Here, we demonstrate that the iiGKBA overcomes the limitations of the conventional GKBA, which fails in this regime, and provides a computationally efficient approach that retains the essential physics of the full Kadanoff-Baym equations. This enables consistent and accurate treatment of time-dependent thermoelectric transport in nanoscale junctions, capturing both transient and steady-state behavior.
@@
In realistic transport setups (see Fig. 1), the form of the system-lead coupling plays a decisive role. It is commonly described by a tunneling rate or line-width function, which characterizes how the states in the central region (e.g., the molecular junction) are broadened due to their interaction with the continuum of states in the leads and determines the rate at which electrons tunnel between the quantum system and the leads. The bandstructure or density of states of the leads is decisive in shaping the energy dependence of the line-width function. While the wide-band limit approximation (WBLA) is widely used for its simplicity [73, 74], it fails to capture these effects. By concentrating on Lorentzian tunneling-rate functions, we present a detailed derivation of the iterated generalized Kadanoff-Baym ansatz and establish how it eliminates unphysical divergences in the standard GKBA, achieving benchmark agreement with full KBE solutions.
Finally, we put the developed iiGKBA scheme to the test by investigating the real-time buildup of thermoelectric energy conversion in a molecular junction under a temperature gradient. This reveals transient efficiency windows, i.e., periods where performance can exceed the steady state, establishing dynamics as essential for understanding nanoscale energy conversion and for guiding the development of future quantum and energy-harvesting technologies.
@@ II.2 Non-equilibrium Green’s function theory
The central object of the non-equilibrium Green’s function theory, in describing the out-of-equilibrium quantum-transport setting of Eq. (1) is the one-electron lesser (<<) Green’s function (GF) [45]
@@
where the labels $r,s$ refer to either the lead or central region states, the creation/annihilation operators are represented in the Heisenberg picture, and the ensemble average $\langle\cdots\rangle$ is taken as a trace over the equilibrium density matrix. This function naturally captures quantum transport processes, as electrons continuously hop onto and off the central molecule, thereby changing the particle number in the central region. The associated one-electron density matrix is obtained as the equal-time limit $\rho(t)\equiv-iG^{<}(t,t)$.
@@
$=\int d\bar{t}\left[\Sigma^{<}(t,\bar{t})G^{\mathrm{A}}(\bar{t},t^{\prime})+\Sigma^{\mathrm{R}}(t,\bar{t})G^{<}(\bar{t},t^{\prime})\right],$ (3)
where $h_{mn}^{\mathrm{HF}}(t)=h_{mn}(t)+V_{mn}^{\mathrm{HF}}(t)$ is the one-electron Hamiltonian including the Hartree-Fock (HF) potential $V_{mn}^{\mathrm{HF}}(t)=\sum_{pq}(v_{mpqn}(t)-v_{mpnq}(t))\rho_{qp}(t)$. The self-energy $\Sigma=\Sigma_{c}+\Sigma_{\mathrm{em}}$ is composed of the correlation and embedding parts. The correlation self-energy can be specified at different levels of approximations: second-order Born [48, 75], $GW$ and $T$-matrix [76, 50], or based on the non-equilibrium dynamical mean-field theory [77, 78, 79]. The embedding self-energy accounts for the system being open to an environment consisting of the leads. The superscripts $\mathrm{A}$’ and $\mathrm{R}$’ denote the advanced and retarded components, respectively.
@@ III.1 Collision integrals
$I_{\alpha}^{\nu}(t)=\int\!dx\bigl[\Sigma_{\alpha}^{\nu,<}(t,x)G^{\mathrm{A}}(x,t)+\Sigma_{\alpha}^{\nu,\mathrm{R}}(t,x)G^{<}(x,t)\bigr].$ (9)
@@ III.2 Reconstruction equations
It is convenient to split the lesser GF into the retarded- and advanced-like ($\mathrm{R}$$\mathrm{A}$) components
@@
$G^{<}_{\mathrm{R}}(t,x)=-G^{\mathrm{R}}(t,x)\rho(x)-\int\!d{y}\,G^{\mathrm{R}}(t,y)\theta(y-x)\\ \times\left\{\int\!d{z}\,\Sigma^{<}(y,z)G^{\mathrm{A}}(z,x)-\Sigma^{\mathrm{R}}(y,z)G^{<}_{\mathrm{A}}(z,x)\right\},$ (11)
@@
$G^{<}_{\mathrm{A}}(x,t)=-\rho(x)G^{\mathrm{A}}(x,t)\\ -\int d{y}\left\{\int\!d{z}\,G^{\mathrm{R}}(x,z)\Sigma^{<}(z,y)+G^{<}_{\mathrm{R}}(x,z)\Sigma^{\mathrm{A}}(z,y)\right\}\\ \times\theta(y-x)G^{\mathrm{A}}(y,t).$ (12)
One obtaines the “standard” GKBA equations by neglecting the memory corrections (given by the integral over $y$). The iterated GKBA (iiGKBA) is obtained by iterating the equations further, i.e., inserting the GKBA expression for $G^{<}_{\mathrm{A}}$ from the second equation into the first, and $G^{<}_{\mathrm{R}}$ from the first equation into the second. In order to efficiently deal with different kinds of convolution integrals that appear in this work, we introduce short-hand notations:
$\bigl[af\cdot b\bigr](t,t^{\prime})$ $=\smallint\!d{x}\,a(t,x)f(x)b(x,t^{\prime}),$ (13)
@@
This allows to rewrite the iiGKBA equations as follows:
@@ III.3 Embedding self-energy components
where $\phi_{\alpha}(t,t^{\prime})\equiv\int_{t^{\prime}}^{t}d{x}V_{\alpha}(x)$ is the accumulated phase due to the applied voltage. With these ingredients, different components of the embedding self-energy can be constructed as Fourier transforms, $\mathcal{F}[a](\tau)=\int\frac{d\omega}{2\pi}e^{-i\omega\tau}a(\omega)$, of the lead tunneling rates [45]:
@@
$\qquad\times\mathcal{F}\left[\omega^{\nu}\Gamma_{\alpha}(\omega)\right](t-t^{\prime})\theta(t-t^{\prime}),$ (17a)
@@
$\qquad\times\mathcal{F}\left[\omega^{\nu}f_{\alpha}(\omega)\Gamma_{\alpha}(\omega)\right](t-t^{\prime}),$ (17b)
@@
$\bar{\eta}_{\alpha\ell}$ $=\begin{cases}i\frac{\gamma_{\alpha}}{2}\Omega_{\alpha}f_{\alpha}(\epsilon_{\alpha}-i\Omega_{\alpha})&\ell=0,\\ -\frac{\eta_{\ell}}{\beta_{\alpha}}\Gamma_{\alpha}\left(\mu_{\alpha}-i\frac{\zeta_{\ell}}{\beta_{\alpha}}\right)&\ell\geq 1;\end{cases}$ (23)
@@
$\bar{\mu}_{\alpha\ell}$ $=\begin{cases}\epsilon_{\alpha}-i\Omega_{\alpha}&\ell=0,\\ \mu_{\alpha}-i\frac{\zeta_{\ell}}{\beta_{\alpha}}&\ell\geq 1.\end{cases}$ (24)
@@ III.4 Working scheme
Figure 2: Derivation of the EOMs for the iiGKBA correlators. Correlators that originate from the self-energy terms in Eqs. (33) are framed in red. Correlators that appear already at the GKBA level are shaded.
@@
$i\frac{d}{dt}G^{\mathrm{A}}(t^{\prime},t)=\delta(t^{\prime}-t)\\ +G^{\mathrm{A}}(t^{\prime},t)h^{\mathrm{HF}}(t)+\bigl[G^{\mathrm{A}}\cdot\Sigma^{\mathrm{A}}\bigr](t^{\prime},t),$ (33b)
@@ IV Benchmark simulations for a quantum-dot system
There are different levels of approximations outlined in Sec. III. Inserting the GKBA iteratively back into the reconstruction equations (11) and (12) resulted in the full set of $14$ EOMs, which is referred to iiGKBA. Neglecting the correction terms $\mathcal{A}$ in the collision integral (29) reduces to the GKBA framework (only shaded correlators in Fig. 2 are considered) with the embedding self-energies described beyond the WBLA, through the Lorentzian line-width function (16). This finite-bandwidth Lorentzian description can be applied systematically to both the collision integrals (29) and the equations of motion for the advanced propagators (33). We refer to this scheme as “GKBA ($\Omega,\ G^{\mathrm{A}}$)”. Alternatively, the retarded/advanced propagators can be approximated at the wide-band approximation level, i.e., the bandwidth $\Omega$ is finite for the embedding self-energies, but the equations of motion (33) include an effective Hamiltonian $h_{\mathrm{eff}}=h^{\mathrm{HF}}-i\gamma/2$ with the frequency-independent line-width matrix ($\gamma=\sum_{\alpha}\gamma_{\alpha}$; see Eq. (16)). We refer to this as “GKBA ($\Omega,\ h_{\mathrm{eff}}$)” with red-framed correlators in Fig. 2 being additionally neglected. The limiting case, $\Omega\to\infty$ is the wide-band limit approximation, simply referred to as “GKBA (WBLA)”: For the charge density and current, we refer to Ref. [70] and for the energy current, an alternative derivation is presented in the Supplemental Material [87]. The ($i$)GKBA calculations were performed using the cheers code [88, 89].
@@
We focus on a thermoelectric transport setup consisting of a non-interacting two-terminal quantum dot with a temperature gradient and a time-dependent chemical potential difference. This model, studied analytically in Ref. [90], can capture complex dynamic thermoelectric behavior, and our benchmark calculations illustrate how different operating regimes impact efficiency, charge and energy-current responses of the system, and the validity range of the developed theoretical formulation in Sec. III.
@@
Figure 3: Time-dependent charge density $\rho$ (a), charge current $J_{L}$ (b), and energy current $J_{L}^{E}$ (c) at the left-lead interface with the quantum dot being weakly coupled to the leads, $\gamma=\varepsilon/10$, and driven by a sudden gate voltage $u(t)=(5\varepsilon/2)\theta(t)$. Inset in panel (c) displays the equilibrium energy current, before the sudden gate voltage is switched on, in terms of the number of poles. Other parameters are $\mu_{L}=\mu_{R}=0$, $\beta_{L}=\beta_{R}=10/\varepsilon$.
@@
Referring back to Eq. (1), we set the quantum dot energy level to $h(t)=\varepsilon/2+u(t)$, where a sudden gate voltage is applied as $u(t)=(5\varepsilon/2)\theta(t)$. All energies are expressed in units of $\varepsilon>0$, defined relative to the average chemical potential. The quantum dot is non-interacting, i.e., $v=0$ in Eq. (1). The chemical potentials of the left and right leads are first set to $\mu_{L}=\mu_{R}=0$ and there is no temperature gradient $\beta_{L}=\beta_{R}=10/\varepsilon$. In all calculations, we set $\Omega_{\alpha}=\Omega$ for all leads $\alpha$. In Fig. 3, the time-dependent response is analyzed in terms of the charge density on the quantum dot $\rho$ and the charge and energy current at the left-lead interface $J_{L}$, $J_{L}^{E}$, when the quantum dot is weakly coupled to the leads $\gamma_{L}+\gamma_{R}=\gamma=\varepsilon/10$. We note that $\gamma$ is a scalar for the single-level quantum-dot system. The time evolutions are started from an initially disconnected system $\rho(t\to-\infty)=0$ and adiabatically ramping the coupling between the quantum dot and the leads. As discussed in Ref. [72], with aligned chemical potentials and no temperature gradients, the equilibrium state should exhibit no currents. The comparison of density and charge current to the analytical result of Ref. [90] (“Kara Slimane et al.” in Fig. 3) is excellent already at the WBLA level. When the leads are described with a finite-width Lorentzian, GKBA ($\Omega=10\varepsilon,\ h_{\mathrm{eff}}$) and GKBA ($\Omega=10\varepsilon,\ G^{\mathrm{A}}$), there is a small change in the equilibrium density $\rho(t=0)$. This has no apparent effect on the charge current. The iiGKBA results follow the same behavior.
@@
For the energy current, Fig. 3(c), the situation changes qualitatively. Although the result of Ref. [90] is obtained by numerical integration, the energy current still converges to a physically expected value for a two-lead setup at equilibrium $J_{L}^{E}(t=0)=0$. The iiGKBA calculation matches this fairly well, although the finite-bandwidth $\Omega=10\varepsilon$ does not precisely correspond to the WBLA. In contrast, for the standard GKBA approach, at different levels of approximation, the corresponding result changes with the number of poles in the expansion (21). The main panels in Fig. 3 include $N_{p}=30$ poles, and the pole count is varied at the inset of panel Fig. 3(c), showing a slow divergence (logarithmic) in the case of WBLA. Interestingly, the GKBA calculations with finite-width Lorentzians do converge and the convergence is relatively fast, around $N_{p}\gtrsim 10$ poles, but they converge to an unphysical value ($\neq 0$). The overall mismatch in the energy currents is still fairly minimal at weak coupling.
@@
Figure 4: Same as Fig. 3 but the quantum dot being strongly coupled to the leads, $\gamma=\varepsilon$.
@@
The iiGKBA scheme accurately resolves the energy currents in the strong coupling regime as well; see Fig. 4. Here, the parameters are otherwise the same as in the weak-coupling calculation (Fig. 3) but the coupling strength is varied so that $\gamma=\varepsilon$. Interestingly, the charge density and current [Fig. 4(a) and Fig. 4(b)] are accurately described by the standard GKBA, but the energy current again diverges for the WBLA case with increasing pole count. We clearly see that when the coupling strength is increased, the standard GKBA at the level of WBLA becomes entirely insufficient. The GKBA calculations with finite-band Lorentzian converge again around $N_{p}\gtrsim 10$ poles, but there is still a noticeable shift in the equilibrium energy current. The iiGKBA corrects this unphysical behavior, but it is worth noting that the finite-bandwidth Lorentzian with $\Omega=10\varepsilon$ description deviates more from the result of Ref. [90], obtained at the WBLA, when the coupling is stronger. Still, it is required to go beyond GKBA to fix the problem of unphysical energy currents in equilibrium.